# Unsourced Random Access: A Recent Paradigm for Massive Connectivity

### v1

@@ Abstract
The sixth generation and beyond communication systems are expected to enable communications of a massive number of machine-type devices. The traffic generated by some of these devices will significantly deviate from those in conventional communication scenarios. For instance, for applications where a massive number of cheap sensors communicate with a base station (BS), the devices will only be sporadically active and there will be no coordination among them or with the BS. For such systems requiring massive random access solutions, a new paradigm called unsourced random access (URA) has been proposed. In URA, all the users employ the same codebook and there is no user identity during the data transmission phase. The destination is only interested in the list of messages being sent from the set of active users. In this survey, we provide a comprehensive overview of existing URA solutions with an emphasis on the state-of-the-art, covering both algorithmic and information-theoretic aspects. Moreover, we provide future research directions and challenges, and describe some potential methods of addressing them.
@@
Multiple access, random access techniques, grant-free, unsourced random access, machine-type communications, Internet of things
@@ I Introduction
Machine Type Communications (MTC) refers to the communication paradigm tailored for the interaction between devices without direct human intervention. It is a key aspect of the Internet of things (IoT) paradigm, enabling various machines, sensors, and devices to exchange information [1, 3, 2]. MTC plays a crucial role in many applications such as smart cities, industrial automation, and healthcare [5, 1, 4]. MTC exhibits distinctions from human-type communication in several aspects. One notable feature is the ability of MTC to handle a large number of devices with sporadic activity for each device [8, 5, 7, 6]. Specifically, MTC is anticipated to accommodate a density of up to 1 Million devices per square kilometer, and these devices typically operate with low computational and storage capabilities, low duty cycles, and small payloads [5, 6].
Multiple access techniques provide the opportunity for different users to share common communication media simultaneously, thus playing a fundamental role in wireless communication systems [9]. Centralized access (referred to as coordinated access) and random access (also known as initial access and uncoordinated access) are two main categories of multiple access technology [6, 8, 11, 10, 4, 12, 9]. Centralized access refers to a communication protocol or technique where access to a shared communication channel is controlled and managed by a central authority, which allocates distinct communication resources to each user and allows them to transfer data. On the contrary, in random access, users send their messages simultaneously as soon as the new data is generated, without the need for explicit coordination with the central unit [11, 10, 13, 4, 14, 12].
Coordinated multiple access stands as the predominant multiple access mode in the fifth generation (5G) and preceding technologies. Several fundamental coordinated access techniques have been introduced from the first generation (1G) to the fourth generation (4G), including time-division multiple access (TDMA), frequency-division multiple access (FDMA), orthogonal frequency-division multiplexing (OFDMA), code-division multiple access (CDMA), and space-division multiple access (SDMA). These techniques belong to the large category of orthogonal multiple access (OMA) schemes. In the context of 5G networks, the utilization of non-orthogonal multiple access (NOMA) improves the capacity region compared to the conventional techniques [15]. To manage the communication of available users in a system, the mentioned techniques require appropriate scheduling protocols, which are efficient only in the case of few active users [16, 17, 12]. However, to accommodate a large number of users in massive MTC (in beyond 5G and the sixth generation (6G) applications), the coordinated protocols incur excessively high access latency and prohibitive signaling overhead [14, 4, 18, 16, 19, 17, 12].
In contrast to the conventional centralized access, where the base station (BS) allocates resources to the users (devices), users transmit their data without any coordination in the random access setup. Removing the need for scheduling results in some benefits, such as reducing the latency and signaling overhead, which makes the random access scheme an interesting solution for massive MTC systems [17, 14, 4, 18, 16, 19]. ALOHA and grant-free multiple access are two common random access schemes [17, 20].
A widely used random access scheme is the ALOHA multiple access protocol, which enables users to transmit data without prior coordination. In the original description given in [21], Abramson explained that in an ALOHA network, a single broadcast channel is used by multiple communication devices, each of which independently sends packets without coordinating with other devices or following a specific schedule. Specifically, a device sends the entire packet at a randomly selected time, and then waits for an acknowledgment for a period of time. If no acknowledgment is received within the timeout period, it is assumed that a collision has occurred with a packet sent by another device. In such cases, the packet is retransmitted after an additional random waiting time to avoid repeated collisions. This process is repeated until a successful transmission and its subsequent confirmation occurs or until the process is terminated by the user console. Slotted ALOHA (SA), a variation of ALOHA, is presented as a method for transmitting packets in designated time slots, which doubles the maximum throughput compared to the pure ALOHA protocol [22]. However, there are still some challenges in SA. For instance, in this protocol, several active users may share the same slot to send their data packets, leading to packets colliding within that slot and the discarding of data [23, 21, 24]. Many studies have focused on improving the effectiveness of SA through various methods such as multiple packet transmission, interference cancellation, and error correction codes [26, 25, 27, 28]. Further analyses of ALOHA-based techniques are provided in [29, 30, 31, 32].
Grant-free random access is another common random access method [5, 12]. In contrast to ALOHA, where users solely transmit data, the grant-free method involves each user transmitting both a preamble and the data sequence [20]. Preamble transmission aims to facilitate collision reduction at the physical layer. Hence, unlike ALOHA, which experiences a significant decline in performance due to collisions, grant-free random access solutions are more resilient to collisions by employing techniques such as user scheduling via a small level of feedback [33], or device-specific patterns such as codebook structures and spreading sequences [17]. So, when there is a large number of users, grant-free random access is a more effective solution than ALOHA. Sourced and unsourced random access schemes are the main categories of grant-free random access. In the first one, both the messages and the identities of users matter to the BS. Therefore, a distinct pilot/preamble is defined for each user, leading to the requirement of a codebook with dimensions proportional to the square of the potential user load. This is impractical in many next-generation wireless network applications with hundreds of millions of potentially connected devices [34, 35].
The so-called unsourced random access (URA), which was introduced by Polyanskiy in [36], by employing a common codebook among users, decouples the user identification and data transmission in the multiple access problems. So, in URA, the receiver is interested in the list of transmitted messages with the decoding done up to a permutation of the messages, and the identity of the users is not important at the first stage. Removing user identity in the URA system enhances its efficiency compared to the sourced random access, enabling it to support a significantly larger number of users [34, 35]. In this paradigm, the BS is connected to millions of cheap devices, a small fraction being active at any given time. The users share all the resources without any priority among them, and they share the same codebook to transmit their payload. Typically, the per-user probability of error (PUPE) is adopted as the performance criterion [36]. The URA paradigm is a promising solution to some next-generation communication applications such as massive machine-to-machine (M2M) communications (one of the main enablers of massive IoT). A block diagram of different multiple access techniques is depicted in Fig. 1.
@@
Following Polyanskiy’s pioneering work, several low-complexity coding solutions are developed in the subsequent literature. In those solutions, by employing successive interference cancellation (SIC) techniques, collisions are resolved up to a certain level depending on user transmit powers. Many techniques, such as random spreading [37], multistage orthogonal pilots [38], and frame slotting [39], have been employed to mitigate the collisions and interference. While a single-antenna receiver in a Gaussian multiple access channel (GMAC) setup is considered in the earlier works, multiple input multiple output (MIMO) receivers, especially massive MIMO, are considered with fading channel in the subsequent literature to support an even greater number of users [37, 38, 39]. Massive MIMO systems [40] help URA by providing high spectral efficiencies, high energy efficiencies, high data rates, and spatial multiplexing gains. This is achieved by generating a substantial number of spatial degrees of freedom (DoF).
Figure 1: Taxonomy of different multiple access techniques.
In the following sections, we provide an overview of the related URA literature, with a focus on GMAC, fading multiple access, and MIMO fading multiple access scenarios. From another aspect, we study the works that put forth coding schemes and those that offer fundamental limit calculations for the URA setup.
The remainder of the survey is organized as follows. In Section II, a general discussion of multiple-access techniques is provided. We start with coordinated multiple-access techniques, both grant-free and grant-based, and find our way to the uncoordinated ones. In Section III, we focus on URA as a new paradigm for uncoordinated multiple-access communications with a massive number of potential devices, enumerate its advantages, and provide some solutions developed in the literature to address the related challenges. Sections IV and V are dedicated to in-depth overviews of the existing URA solutions, with Section IV focusing on URA over GMAC and Section V on URA over fading channels. Regarding communication models over practical wireless channels, we first investigate the fading works with a single antenna receiver and describe the general approaches; then, we discuss those with a MIMO assumption. Finally, we conclude the survey in Section VI with remarks on URA solutions and provide future research directions.
The acronyms used in this paper are alphabetically summarized in Table I.
Table I: List of acronyms
@@
CB-MLD covariance-based maximum likelihood detection
@@
CDMA code division multiple access
@@
FDMA frequency division multiple access
@@
GMAC Gaussian multiple access channel
@@
IDMA interleave-division multiple access
@@
ML maximum likelihood
@@
NOMA non-orthogonal multiple access
@@
ODMA on-off division multiple access
@@
SCMA sparse code multiple access
SDMA space division multiple access
@@
TDMA time division multiple access
@@
URA unsourced random access
@@ II Background: Multiple Access
II-A Coordinated Grant-Based Multiple Access
@@
Multiple access refers to the scenario in which more than one user simultaneously shares the system resources. One way to serve multiple users is to assign dedicated resources (e.g., time, frequency, code, or their combination) to them by a central coordination unit through a prior handshaking procedure. The conventional techniques in this framework are TDMA, FDMA, CDMA, OFDMA, and SDMA, which are all grant-based techniques. In TDMA and FDMA, non-overlapping time slots or frequency sub-channels are assigned to the different users, respectively, and each user utilizes its allocated time frame or frequency band for transmission. On the other hand, in CDMA, all the users can simultaneously utilize all the time-frequency resources through different (nearly) orthogonal code sequences coupled with a low-complexity decorrelation-based receiver. OFDMA provides frequency division multiplexing by assigning different sets of subchannels to the individual users in an orthogonal frequency division multiplexing (OFDM) system, while SDMA relies on the principle of creating (nearly) orthogonal spatial channels for different users to minimize the inter-user interference by employing beamforming and spatial multiplexing.
@@
All of the approaches described in the previous paragraph are in the category of OMA, where the number of users is strictly limited due to the orthogonal resource assignment. In order to increase the number of supported users, NOMA techniques are investigated in the literature, where multiple users can utilize the same resources to transmit their data, allowing the system to support more users. However, this imposes inter-user interference on the system, and more sophisticated interference-canceling receivers are needed with an increased complexity compared to OMA.
@@
The NOMA schemes are divided into two main categories: power-domain NOMA and code-domain NOMA. In power-domain NOMA [41], the users employ different power levels to transmit their data using the same resources, and the receiver exploits the received power differences for detection through SIC. On the other hand, code-domain NOMA is similar to classical CDMA in the sense that the users employ user-specific code sequences to transmit their data. The code sequences are sparse or non-orthogonal sequences with low cross-correlations [42]. Here, the receiver can employ message-passing algorithms utilizing the code structure to differentiate the users. Some examples of code-domain NOMA schemes in the literature are low-density spreading code division multiple access [43], low-density spreading aided orthogonal frequency division multiplexing [44], sparse code multiple access (SCMA) [45], and multiuser shared access [46].
II-B Grant-Free Multiple Access
@@
All of the schemes described in the previous subsection are grant-based. That is, they require scheduling grants and coordination with the BS before transmission, and it is assumed that the number of users and their activity patterns are available at the BS. However, when the number of users becomes large, they become infeasible due to the excessive delays and signaling overhead. For example, the typical number of users per signal dimension in the time-frequency domain (per degree of freedom) is between 1.5-3 according to 3GPP studies, as the system performance significantly degrades beyond that threshold [47].
Grant-free multiple access is a well-known candidate to enable communication of a large number of users where the users transmit their data by employing the shared system resources without any prior grant from the BS, which reduces the latency and signaling overhead. Nevertheless, the users can have unique signatures/preambles or they can utilize different codebooks pre-configured by the BS, which means that there is user identity during the data transmission phase, and the receiver can utilize this information for activity detection or user separation. In this line, grant-free NOMA is the prominent solution for future communication systems due to its low latency, improved spectral efficiency, and potential to support massive connectivity. Note that the main difference between grant-based and grant-free NOMA is that, the BS does not have information on the set of active users in the latter, which brings significant challenges requiring different solutions such as blind activity detection and data recovery, channel estimation, and synchronization with minimal overhead [12].
Among the two main categories of NOMA, the power-domain NOMA is hard to extend to the grant-free scenario as the successful recovery of the transmitted packets at the receiver heavily depends on the power difference of the users, which may not be maintained due to the grant-free nature of the transmissions. On the other hand, signature-based schemes in code-domain NOMA can be adapted to the grant-free case with proper activity detection before the data detection or joint activity detection and data recovery. For instance, a grant-free version of SCMA is proposed in [48]. Furthermore, the inherent sparsity of the user activity in grant-free multiple access is utilized in [49, 50, 51, 52] through compressive sensing algorithms for multiuser detection.
II-C Uncoordinated Multiple Access
Another candidate solution to address the drawbacks of coordinated multiple access is uncoordinated multiple access, where the users independently exploit the system resources to transmit their data using the same transmission protocol without any coordination with the BS. This scheme is particularly useful when there are stringent latency requirements or when the user activity is random/sporadic, which renders assigning dedicated system resources ineffective.
The first approach for uncoordinated multiple access is the ALOHA protocol [23] proposed in the 1970s, where the users transmit their packets whenever they have data to send. The data packets are assumed to be lost in the case of a collision and they should be re-transmitted. ALOHA is simple; however, it is inefficient as the system throughput is $T_{e}=Le^{-2L}$ with a maximum of $1/{2e}\approx 18\%$ where $T_{e}$ is the system throughput and $L$ is the system load defined as the ratio of the number of transmitted packets to the number of slots. An improvement to the pure version of ALOHA is slotted ALOHA, in which it is assumed that the time is divided into slots and the users can start the transmission only at the beginning of the time slots. Thus, there are no partial collisions, and a packet is assumed as successfully received if there are no other transmissions in the same slot. In this case, the throughput can be calculated as $T_{e}=Le^{-L}$, with a maximum of $1/{e}\approx 37\%$. The transmission structures in pure ALOHA and slotted ALOHA are illustrated in Fig. 2, and their throughput with respect to the system load is depicted in Fig. 3.

### v2

# Unsourced Random Access: A Comprehensive Survey

@@ Abstract
Multiple access communication systems enable numerous users to share common communication resources simultaneously, playing a crucial role in wireless networks. With the emergence of the sixth generation (6G) and beyond communication systems, supporting massive machine-type communications with sporadic activity patterns is expected to become a critical challenge. Unsourced random access (URA) has emerged as a promising paradigm to address this challenge by decoupling user identification from data transmission through the use of a common codebook. This survey provides a comprehensive overview of URA solutions, covering both theoretical foundations and practical implementations. We present a systematic classification of URA solutions across three main channel models: Gaussian multiple access channels (GMACs), single-antenna fading, and multiple-input multiple-output (MIMO) fading channels. For each category, we analyze and compare state-of-the-art solutions in terms of performance, complexity, and practical feasibility. Additionally, we discuss critical challenges such as interference management, computational complexity, and synchronization issues. The survey concludes with promising future research directions and potential methods to address existing limitations, providing a roadmap for researchers and practitioners in this rapidly evolving field.
@@
Multiple access, RA techniques, grant-free, unsourced RA, machine-type communications, Internet of things
@@ I Introduction
The rapid proliferation of Internet of Things (IoT) devices is driving a fundamental shift in wireless communication systems. Current estimates project over 15 billion connected IoT devices by 2023, with the expected increase to a staggering number of over 38 billion devices by 2029 [1]. One example of such systems is machine-type communications (MTC) (also known as machine-to-machine (M2M) communications), which, as a key aspect of the IoT paradigm, enables various machines, sensors, and devices to exchange information without direct human intervention. MTC plays a crucial role in many applications ranging from industrial automation and smart cities to healthcare monitoring and environmental sensing [2]. Specifically, massive MTC is anticipated to accommodate a density of up to one million devices (per square kilometer), which typically operate with low computational and storage capabilities, low duty cycles, and small payloads [3]. This massive scale presents unprecedented challenges for wireless access, as traditional communication paradigms were designed for a relatively small number of high-data-rate users rather than an enormous number of devices with sporadic, low-rate transmissions.
Multiple access (MA) techniques, which enable different users to share a common communication medium simultaneously, have been fundamental to wireless communication systems [4]. MA technologies can be broadly categorized into centralized access and random access (RA). In centralized access (also known as coordinated access), a central authority controls and manages access to the shared communication channel by allocating distinct resources to each user for data transmission. This approach has been the predominant MA mode in the fifth generation (5G) and previous cellular technologies. However, centralized coordination becomes impractical for massive IoT scenarios due to prohibitive signaling overhead and coordination complexity. In contrast, in RA (also known as uncoordinated access), users transmit their messages simultaneously as soon as new data is generated, without requiring explicit coordination with the central unit. This approach reduces both latency and signaling overhead. These advantages make RA solutions particularly suitable for next-generation (6G and beyond) communication networks, especially when serving a large number of users [5, 6].
Figure 1: Taxonomy of different MA techniques.
Traditional RA solutions like ALOHA [7] and its variants suffer from excessive collisions as the number of users grows. More sophisticated grant-free RA schemes [8] employ unique signatures or preambles for each potential user to facilitate collision reduction at the physical layer. Grant-free RA solutions can accommodate more users compared to ALOHA, as they can employ user scheduling via a limited amount of feedback [9], or device-specific patterns such as codebook structures and spreading sequences [10]. However, the distinct pilot/preamble requirement for each potential user becomes infeasible as the number of connected devices reaches hundreds of millions. On the other hand, in many of the emerging applications, such as such as massive MTC and IoT, devices become active only sporadically. Specifically, massive MTC is anticipated to accommodate a density of up to one million devices (per square kilometer) with low computational and storage capabilities, low duty cycles, and small payloads [3].
@@
To mitigate the challenge of identifying and decoding messages from such a large number of potential devices, Polyanskiy introduced unsourced random access (URA) paradigm [11]. URA fundamentally reimagines the massive access problem by decoupling user identification from data transmission. Instead of assigning unique preambles or codebooks to each potential user employ a common codebook, and the receiver focuses solely on recovering the list of transmitted messages. This paradigm shift drastically reduces system complexity and enables practical implementations for truly massive numbers of devices. Following Polyanskiy’s pioneering work, several low-complexity coding solutions have been developed in the subsequent literature, in which, by employing successive interference cancellation (SIC) techniques, collisions are resolved up to a certain level depending on user transmit powers. Many techniques, such as random spreading [12], multistage orthogonal pilots [13], and time frame slotting [14], have been employed to mitigate the collisions and interference. While a single-antenna receiver in a Gaussian MA channel (GMAC) setup is considered in the earlier works, the literature is shifted to fading channels with multiple-antenna receivers, especially massive multiple-input multiple-output (MIMO), to support more practical channels and an even greater number of users. A block diagram of different MA techniques is depicted in Fig. 1 alongside the three main set-ups in URA research, namely, GMAC, and single-antenna and MIMO fading channels. The figure also shows the way this survey is structured.
Although close to ten years have passed since Polyanskiy’s initial work, and unsourced random access has been studied in many papers, there have been limited efforts to review the existing schemes on URA. The first attempt is a survey on massive access [15], the authors review both many-user MA, where each user is assigned an individual codebook (i.e., the coordinated approach), and URA, which they call uncoordinated massive access. However, since URA was in its early stages at the time of this survey, it mostly covers the earlier works of GMAC URA. In [16], the focus is again on GMAC URA, and the authors review the design paradigms and existing coding schemes for URA over GMAC while fading channels and MIMO URA are presented as open problems. In a less URA-oriented survey [17], an overview of compressive sensing-based grant-free massive access is covered, where some compressed sensing-based URA schemes are also included. In a recent survey [18], through an overview of the existing URA schemes studying mostly GMAC, the authors present URA as a suitable candidate for next-generation communication systems and show that it can support the 3rd Generation Partnership Project (3GPP) standardization.
While the existing surveys focus on the GMAC setup, this survey also covers the fading URA solutions, which comprise the majority of works in the URA literature. More specifically, we conduct a comprehensive review of the existing URA solutions in the literature in its three main setups, namely, URA over GMAC and single-antenna and MIMO fading channels, accompanied by a discussion on the main challenges with possible solutions and many promising future research directions. We study works that put forth coding schemes as well as those that offer fundamental information-theoretical limit calculations for the URA setup. To the best of our knowledge, this is the first comprehensive survey on URA, reviewing this crucial emerging paradigm for 6G and beyond communications in all aspects. A summary of the existing surveys and our contribution can be found in Table I.
Table I: Summary of the surveys on URA.
Paper Scope
Wu et al., 2020 [15] Review of many-user MAC and URA with a focus on GMAC.
Li et al., 2022 [16] Review of basics of URA and coding schemes on URA over GMAC.
Gao et al., 2024 [17] Review of grant-free massive access with a focus on compressed sensing-based schemes.
Liva et al., 2024 [18] Review of URA over GMAC and showing that it can support 3GPP standardization.
This survey A comprehensive review of the URA literature, covering all existing solutions in GMAC and single-antenna and MIMO fading channels.
Table II: List of acronyms
@@
CB-MLD covariance-based MLD
@@
CDMA code division MA
@@
FDMA frequency division MA
@@
FL federated learning
GMAC Gaussian MA channel
@@
IDMA interleave-division MA
@@
ML machine learning
MLD maximum likelihood detection
@@
NOMA non-orthogonal MA
@@
ODMA on-off division MA
@@
SCMA sparse code MA
SDMA space division MA
@@
TDMA time division MA
@@
URA unsourced RA
The remainder of the survey is organized as follows. In Section II, a general discussion of MA techniques is provided. We start with coordinated MA techniques and find our way to random access. In Section III, we focus on URA as a new paradigm for uncoordinated MA communications with a massive number of potential devices, enumerate its advantages, and hint at some solutions developed in the literature to address the related challenges. Sections IV and V are dedicated to in-depth overviews of the existing URA solutions, with Section IV focusing on URA over GMAC and Section V on URA over fading channels. Regarding communication models over practical wireless channels, we first investigate the fading works with a single antenna receiver and describe the general approaches; then, we discuss those with a MIMO assumption. Finally, we conclude the survey in Section VI with remarks on URA solutions and provide future research directions.
The acronyms used in this paper are alphabetically summarized in Table II.
@@ II Background: Multiple Access
Multiple access refers to the scenario in which two or more users simultaneously share the system resources. In this section, we provide a concise overview of the main MA approaches and enumerate their pros and cons in order to better understand the position of URA solutions in next-generation communication systems.
II-A Coordinated Access
@@
One way to serve multiple users is to assign dedicated resources (e.g., time, frequency, code, or their combinations) to them by a central coordination unit through a prior handshaking procedure. The conventional techniques in this framework are time-division MA (TDMA), frequency-division MA (FDMA), orthogonal frequency-division MA (OFDMA), code-division MA (CDMA), and space-division MA (SDMA). In TDMA and FDMA, non-overlapping time slots and frequency sub-channels are assigned to the different users, respectively, and each user utilizes its allocated time frame or frequency band for transmission [4]. On the other hand, in CDMA, all the users can simultaneously utilize all the time-frequency resources through different (nearly) orthogonal code sequences coupled with a low-complexity decorrelation-based receiver. OFDMA provides frequency division multiplexing by assigning different sets of subchannels to the individual users in an orthogonal frequency division multiplexing (OFDM) system, while SDMA relies on the principle of creating (nearly) orthogonal spatial channels for different users to minimize the inter-user interference by employing beamforming and spatial multiplexing [19].
@@
All of the approaches described in the previous paragraph are in the category of OMA, where the number of users is strictly limited due to the orthogonal resource assignment. In order to increase the number of supported users, non-orthogonal MA (NOMA) techniques are investigated in the literature, where multiple users can utilize the same resources to transmit their data, allowing the system to support more users. However, this imposes inter-user interference on the system, and more sophisticated interference-canceling receivers are needed with an increased complexity compared to OMA.
@@
NOMA schemes are divided into two main categories: power-domain NOMA and code-domain NOMA. In power-domain NOMA [20], the users employ different power levels to transmit their data using the same resources, and the receiver exploits the received power differences for detection through successive interference cancellation (SIC). On the other hand, code-domain NOMA is similar to classical CDMA in the sense that the users employ user-specific code sequences to transmit their data. The code sequences are sparse or non-orthogonal sequences with low cross-correlations [21]. Here, the receiver can employ message-passing algorithms utilizing the code structure to differentiate the users. Some examples of code-domain NOMA schemes in the literature are low-density spreading code division MA [22], low-density spreading aided orthogonal frequency division multiplexing [23], sparse code MA (SCMA) [24], and multiuser shared access [25].
II-B Random Access
@@
All of the schemes described in the previous subsection are coordinated. That is, they require scheduling grants and coordination with the BS before transmission, and it is assumed that the number of users and their activity patterns are available at the BS. However, when the number of users becomes large, they become infeasible due to the excessive delays and signaling overhead. For example, the typical number of users per signal dimension in the time-frequency domain (per degree of freedom) is between 1.5-3 according to 3GPP studies, as the system performance significantly degrades beyond that threshold [15]. The common solution addressing these drawbacks is random access, where the users perform their transmission without any coordination by the BS whenever they have data. ALOHA and grant-free random access are the two common RA approaches in the literature.
II-B1 ALOHA & Variants
An early RA approach is the ALOHA protocol [7] (also known as pure ALOHA) proposed in the 1970s, where a device sends the entire packet at a randomly selected time. If no acknowledgment is received within a certain period of time, it is assumed that a collision has occurred with a packet sent by another device, and the packet is retransmitted after an additional random waiting time to avoid repeated collisions. This process is repeated until a successful transmission and its subsequent confirmation occur or until the process is terminated by the user console. Slotted ALOHA (SA) is a variant of ALOHA, where the packets are transmitted in designated time slots, doubling the maximum throughput compared to the pure ALOHA [26]. However, collision is still a major challenge in SA, as several active users may share the same slot to send their data packets, leading to packets colliding within that slot and the discarding of data. The transmission structures in pure ALOHA and slotted ALOHA are illustrated in Fig. 2.